OpenStudy (anonymous):

Can someone please factor out for me? I am absolutely stuck. Is it even possible?? 4tan^2x - 1 = tan^x

mathslover (mathslover):

Is it : \(4 \tan ^{2x} - 1 = \tan^x \) or \(4 \tan ^2 x - 1 = \tan ^x \)

OpenStudy (anonymous):

I'd reckon its \[4\tan ^{2}x - 1 = \tan x\] since tan needs a variable :S

OpenStudy (anonymous):

Yes, that's the correct equation. Do any of you know what the factored form is?

OpenStudy (anonymous):

Try replacing tan x with u. \[4u ^{2} - 1 = u\] Can you factor this?

OpenStudy (anonymous):

no, not using integers anyway

OpenStudy (anonymous):

So the final answer would be no solution?

OpenStudy (anonymous):

what is the original question?

OpenStudy (anonymous):

I believe you can apply the quadratic formula or complete the square, correct me if I'm wrong. 4u^2 - u - 1 = 0 4 (u^2 - 1/4u) - 1 = 0 4(u^2 - 1/4u + 1/64) - 1 - 1/16 = 0 4(u - 1/8)^2 - 17/16 = 0 Then you can substitute tan x = u back in: 4 (tan x - 1/8)^2 - 17/16 = 0 And this should be the solution.

OpenStudy (anonymous):

4tan^2 x−1=tanx Factor into two parenthesis. I got as far as 4tan^2 x - tanx - 1 = 0

OpenStudy (anonymous):

Okay. Thanks! @DoritosHero

OpenStudy (anonymous):

No problem :D